Circular No. 5798

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA AQUILAE 1993 R. H. McNaught, Anglo-Australian Observatory, reports the following

precise position for the nova from a film taken with the Uppsala Southern Schmidt on May 17.6 UT: α = 19^{h}10^{m}34^{s}.69, δ = +1^{o}29′14′′.2 (equinox 1950.0; uncertainty 0′′.3 in each coordinate). Using the same PPM catalogue stars used for the measurement, the nearest star to the nova on a U.K. Schmidt R plate taken 1987 May 24 is at mag 19 and has end figures 34^{s}.56, 14′′.9; this star appears on the Palomar Sky Survey in the same position, indicating that proper motion is not responsible for the position discrepancy. The star shows neutral colors and there is no evidence of variability on the Palomar prints or on 12 U.K. Schmidt exposures in J, R, and I taken during 1979 to 1988. This suggests that the pre-nova was fainter than mag 22.

Further visual magnitude estimates: May 16.55 UT, 8.0 (P. Williams, Heathcote, N.S.W., Australia); 17.99, 8.0 (H. Dahle, Oslo, Norway); 19.00, 7.7 (P. Rapavý, Rimavská Sobota, Slovakia).

SUPERNOVA 1993M IN ANONYMOUS GALAXY E. Cappellaro, European Southern Observatory, reports: “On May

17.2 UT, I obtained a good S/N spectrogram (range 370-900 nm, resolution 1.5 nm) of SN 1993M (cf. IAUC 5783) using the 3.6-m telescope (+ EFOSC multi-mode instrument) at La Silla. Once a redshift of 0.078 is adopted, the spectrum is identical to that of the type-Ia SN 1989B (Barbon et al. 1990, A.Ap. 273, 79) obtained 11 days after maximum, but the Si II 635.5-nm absorption (measured at 671 nm) is slightly less pronounced. Preliminary photometry obtained on May 16.2 at the 1.5-m Danish telescope gives: B = 21.4, B–V = +0.6, V–R = +0.3.”

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Total visual magnitude estimates (cf. IAUC 5766): Apr. 17.26 UT,

13.3 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); 17.85, 13.2 (H. Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + CCD + V filter); 24.06, 13.5 (M. Tombelli, Montelupo, Italy, 0.40-m reflector); May 17.94, 14.2 (Mikuž).

1993 May 19 Daniel W. E. Green